OpenStudy (anonymous):

Lowering the temperature _____ most chemical reactions. (1 point) slows down speeds up does not affect stops

sam (.sam.):

Based on kinetic theory, \[\frac{1}{2}m<c^2>=\frac{3}{2}k_BT \\ \\ m<c^2>=3k_BT\] You can see that \(<c^2> ∝ T\), The mean squared speed is directly proportional to temperature, therefore, when you lowered the temperature, the mean speed of the molecules are much slower, resulting in slower reactions. It doesn't stop the reaction, it just slows down. (Note: You can also recall that molecules require certain amount of speed and orientation in order to form bonds.)